OpenStudy (anonymous):

find two positive real numbers whose prodcut is a maximum. 1) The sum is 10. 2) Sum of first and twice the second is 24. Please explain not just tell answer. Thanks

OpenStudy (mathmate):

For integers, you only need to tabulate and find the maximum (5*5>4*6>3*7...) For real numbers, and to solve it correctly, you need to use calculus, or the properties of parabolas. Which of these are you working on?

OpenStudy (anonymous):

I can only assume where working on properties of parabolas, because where dealing with parabola graphs now but i missed a whole week of class and just collected the homework.

OpenStudy (mathmate):

If you're working on parabolas, you can try to formulate the problem in the form of a quadratic equation, and use completing squares to find the vertex, which is where the maximum or minimum is.

OpenStudy (anonymous):

Could you tell me how to go about doing so to solve it anyway..... I have the answers for both i just don't quite understand The sum is 10 one.. The answer is 55,55.

OpenStudy (mathmate):

I will solve the first one. You can try the second one in a similar way. The canonical form of a quadratic equation is y=a(x-h)^2+k where a is a constant, (h,k) are the coordinates of the vertex Let x be one of the numbers, then 10-x is the other (sum=10). The product of the two numbers is therefore: y=x(10-x) =10x-x^2 Completing squares =-1(x^2-10x+25) + 25 =-(x-5)^2+25 so the vertex is at (5,25), which means that the numbers are x=5, and x=10-5=5 (both 5). For the second problem, the sum of one number and twice the other is 24, so one number is x, and the other is 24-2x.

OpenStudy (anonymous):

In my textbook it says the answers 55,55 as the two numbers that's the reason I am confused/ I'm a lost cause for this though right now though I appreciate you trying to help me so I award you. :]